# Provable Emergence of Deep Neural Collapse and Low-Rank Bias in $L^{2}$-Regularized Nonlinear Networks

### v2

@@ Abstract
Recent work in deep learning has shown strong empirical and theoretical evidence of an implicit low-rank bias: weight matrices in deep networks tend to be approximately low-rank. Moreover, removing relatively small singular values during training, or from available trained models, may significantly reduce model size while maintaining or even improving model performance. However, the majority of the theoretical investigations around low-rank bias in neural networks deal with oversimplified models, often not taking into account the impact of nonlinearity. In this work, we first of all quantify a link between the phenomenon of deep neural collapse and the emergence of low-rank weight matrices for a general class of feedforward networks with nonlinear activation. In addition, for the general class of nonlinear feedforward and residual networks, we prove the global optimality of deep neural collapsed configurations and the practical absence of a loss barrier between interpolating minima and globally optimal points, offering a possible explanation for its common occurrence. As a byproduct, our theory also allows us to forecast the final global structure of singular values before training. Our theoretical findings are supported by a range of experimental evaluations illustrating the phenomenon.
1 Introduction and related work
@@
Recent years have seen growing theoretical and experimental evidence of low-rank properties of deep neural networks. For example, it has been noted that a large portion of singular values from linear layers in different models can be removed almost without affecting performance, both for feedforward and residual networks Schotthofer_low_rank ; Ydelbayev_low_rank . Analogous results have been shown on feedforward and convolutional networks for image classification Li_basis_CNN ; wang2021pufferfish ; khodak2021initialization ; Schotthofer_low_rank as well as on Recurrent Neural Networks (RNNs) trained to learn neural dynamics in neuroscience tasks schuessler2020interplay ; pellegrino2023low .
@@
Alongside the wealth of experimental evidence of low-rank properties of neural networks, several recent works have investigated the low-rank bias phenomenon from the theoretical point of view, providing evidence of decaying singular values in deep linear architectures arora2019implicit ; feng2022rank ; huh2022low and of low-rank properties of networks’ Hessian maps singh2021analytic , analyzing low-rank geometric properties of training gradient flow dynamics for deep linear networks bah2022learning , showing that low-rank bias is related to weight decay galanti2023characterizing ; sukenik2024neural ; kuzborskij2025lowrankbiasweightdecay , and to training invariances le2022training ; ziyin2025parametersymmetrybreakingrestoration . In particular, the authors relate the rank of the weight matrices during SGD to the batch size adopted during the training of the network. However, based only on the batch size, the results from galanti2023characterizing become uninformative when the batch size is comparable to the number of parameters of the neural network.
@@
Together with this, another related phenomenon often observed during the training of classifiers is the so-called “neural collapse” papyan2020prevalence ; zhu2021geometric , which is often described using four properties named NC1–4. Different works have been proving the optimality of neural collapse in different simplified settings zhu2021geometric ; seleznova2023neural ; jacot2025wide , and different extensions have been proposed for its deep version. In particular, we are interested in the first property of Deep Neural Collapse ($\operatorname{\mathrm{DNC1}}$), which essentially states that the within-class variability of intermediate representations goes to zero even for layers in the middle of a deep neural network rangamani2023feature ; sukenik2024neural ; sukenik2024deep (see Definition˜3.5).
@@
In this work, we show a connection between the phenomenon of deep neural collapse and low-rank implicit bias in general nonlinear networks. The main contribution is theoretical and in particular, our main results, which hold in the presence of weight decay, can be summarized as follows:
@@
  • in Section˜3 we prove that within-class variability of hidden layers during training (including the initial dataset) has a quantifiable effect on the numerical rank of the network parameters as the latter decays with the total cluster variation of the hidden representations;
@@
  • in Section˜4 we show that zero total cluster variation for all intermediate layers is optimal in a fairly general setting, both for feedforward and residual networks (Theorem˜4.2). Moreover, in Theorem˜4.3, we prove that for (almost) every interpolating initial condition, there exists a loss-decreasing trajectory connecting the initial condition to a collapsed configuration, thereby offering an explanation to the consistent empirical observation of deep neural collapse rangamani2023feature .
@@
Overall, the combination of the results above proves that in the presence of a small weight decay regularization, zero within-class variability for all intermediate representations is globally optimal under some natural constraints, therefore showing that global minima in parameter space are low-rank matrices. Moreover, convergence to these highly structured solutions may be regarded as universal in the sense that, in the presence of weight decay there is essentially no loss barrier between interpolating features to deep-neural collapsed ones.
The key points of our investigation are the following observations:
  • The stationary points of weight decay-regularized losses $\mathcal{L}_{\lambda}(\mathit{\Theta},x)=\mathcal{L}_{0}(\mathit{\Theta},x)+\lambda\|\mathit{\Theta}\|^{2}$ are solutions to nonlinear eigenvalue equations
    $\nabla_{W_{l}}\mathcal{L}_{0}(\mathit{\Theta},x)=-\lambda W_{l},$
    where $W_{l}$ is the weight matrix of the $l$-th layer. This eigenvalue equation allows us to transfer properties from the gradient to the parameters and vice versa. A similar connection between parameter ranks and weight decay has recently been used in galanti2023characterizing to address the implicit bias problem from a different perspective.
  • The rank of the gradient of the nonregularized loss $\mathcal{L}$ evaluated on $k$ training examples is controlled by $k$, namely $\operatorname{rank}\nabla_{W_{l}}\mathcal{L}_{0}(\mathit{\Theta},x)\leq k$.
  • For a sequence of layers with weight-decay, it is possible to reparameterize the representation cost as a function of the hidden features, as done in jacot2022featurelearningl2regularizeddnns . This setting is not restrictive since the stationary points of the reparametrized cost are in bijection with the original ones jacot2022featurelearningl2regularizeddnns .
  • It is possible to study exactly the global minima of the representation cost $\mathscr{L}(\mathit{\Theta})=\sum_{l=2}^{L}\|W_{l}\|^{2}$ constrained to the set of interpolating configurations, proving that all intermediate features satisfy $\operatorname{\mathrm{DNC1}}$. Thanks to this last result, it is also possible to study the structure of the interpolating minima of $\sum_{l=1}^{L}\|W_{l}\|^{2}$ as a perturbative limit of the previous one when the norm of the input data is much bigger than the norm of the target matrix, i.e., $\|X\|\gg\|Y\|$ (Proposition˜A.2), or for general data matrices in the large depth regime (Remark˜B.6).
  • In this setting, it is also possible to study mode connectivity of the loss landscape under a set of natural constraints $\mathcal{S}$. In particular, we prove that for (almost) every $\mathit{\Theta}(0)$ such that $\mathcal{L}_{0}(\mathit{\Theta}(0),X)=0$ we have:
    $\inf_{\gamma\in\Gamma}\sup_{t\in[0,1]}\mathscr{L}(\gamma(t))\leq\mathscr{L}(\mathit{\Theta}(0))$
    where $\Gamma:=\{\gamma\in\mathcal{C}^{0}([0,1];\mathcal{S})\,|\,:\gamma(0)=\mathit{\Theta}(0),\gamma(1)\,\text{satisfies}\,\,\operatorname{\mathrm{DNC1}}\,\forall l=1,\dots,L\}$.
Together, these results show that there is essentially no loss barrier between zero loss points and global minima of $\mathscr{L}$, which we prove are all deep-neural collapsed. This will be proven constructively, showing the existence of an energy decreasing path with the desired property. Being this result decoupled from the particular dynamics during training, it might be a step in explaining the universality in the emergence of (deep) neural collapse. Moreover, we highlight that, despite the convexity of $\mathscr{L}$, the constraint $\mathcal{L}_{0}(\mathit{\Theta};X)=0$ is highly nonconvex, making this last result nontrivial.
@@ 2 Setting and preliminaries
to learn from the given data and labels. In equation (1), $\sigma_{l}$ are scalar nonlinear activation functions acting entrywise and $\mathit{\Theta}$ denotes all trainable parameters $\mathit{\Theta}=(W_{1},\dots,W_{L},b_{1},\dots,b_{L})$. We write
@@
to denote the training parameters and concatenation of layers from $l_{1}+1$ to $l_{2}$, respectively. We adopt the convention that, when we do not specify the lower index then it starts from zero, and when we omit the top one it ends at the overall depth $L$, namely, $\mathit{\Theta}^{l_{2}}:=\mathit{\Theta}_{0}^{l_{2}}$, $f^{l_{2}}:=f^{l_{2}}_{0}$ and $\mathit{\Theta}_{l_{1}}:=\mathit{\Theta}_{l_{1}}^{L}$, $f_{l_{1}}:=f^{L}_{l_{1}}$. Notice that, with this notation, $\mathit{\Theta}=\mathit{\Theta}_{0}=\mathit{\Theta}^{L}=\mathit{\Theta}_{0}^{L}$, $f=f_{0}=f^{L}=f_{0}^{L}$ and $\mathit{\Theta}^{0}$ is empty. Finally, for the sake of brevity, we write $\mathcal{Z}^{l}:=\{z^{l}_{i}\}_{x_{i}\in\mathcal{X}}$ to denote the set of outputs of the $l$-th layer, where $z^{l}_{i}=f^{l}(\mathit{\Theta}^{l},x_{i})$ is defined as in (1) with input data $x_{i}$, and $\mathcal{Z}^{0}=\mathcal{X}$.
@@
3 Neural Rank Collapse: a link between low-rank and deep-neural collapse
@@
Equation (4) corresponds to an “eigenvalue equation” for the nonlinear operator $\nabla\operatorname{\mathcal{L}}_{0}$. In other words, for a fixed weight decay parameter $\lambda$, training the neural network corresponds to finding a nonlinear eigenvector $\mathit{\Theta}^{*}$ having $\lambda$ as eigenvalue. In addition, equation (4) immediately implies that whenever the operators $\nabla_{W_{l}}\operatorname{\mathcal{L}}_{0}$ have low rank (i.e., they take values in some low-rank space), then also any weight matrix $W^{*}$ composing the eigenvector $\mathit{\Theta}^{*}$ has necessarily low rank. In the following, we investigate this relationship more in depth, providing sufficient conditions to have approximately low-rank gradients.
@@
Let $f(\mathit{\Theta},x)$ be a feedforward neural network as in (1) and $\operatorname{\ell}:\mathbb{R}^{c}\times\mathbb{R}^{c}\to\mathbb{R}$ be a differentiable loss function as in (3). Then, for any $\{(x_{i},y_{i})\}_{i=1}^{K}\subseteq\mathbb{R}^{d}\times\mathbb{R}^{c}$, we have $\operatorname{rank}\Big{(}\operatorname{\nabla}_{W_{l}}\sum_{i=1}^{K}\operatorname{\ell}\big{(}f(\mathit{\Theta},x_{i}),y_{i}\big{)}\Big{)}\leq K$.
@@
In addition, Proposition˜3.1 immediately yields an example where gradients have necessarily low rank. Consider the case of degenerate data, i.e., assume that there exist $K$ pairs $\{(\bar{x}_{k},\bar{y}_{k})\}_{k=1}^{K}$ such that any data point $(x_{i},y_{i})$ is equal to some $(\bar{x}_{k},\bar{y}_{k})$. Then, by Proposition˜3.1, $\operatorname{\nabla}_{W_{l}}\sum_{i}\operatorname{\ell}(\mathit{\Theta},x_{i},y_{i})$ has rank at most $K$. But the right-hand side of (4) has rank at most $K$, hence any critical point of the loss function $\operatorname{\mathcal{L}}_{\lambda}$ necessarily corresponds to a neural network having all layers of rank at most $K$. The same argument can be extended to the situation of total intermediate neural collapse Poggio_hidden_nc , where the intermediate hidden set $\mathcal{Z}^{l}$ consists of at most $K$ distinct points. Indeed, assume that for a stationary point $\mathit{\Theta}^{*}$ there exists a layer $l$ and points $\{(\bar{z}_{k}^{l},\bar{y}_{k})\}_{k=1}^{K}$ such that any pair $(z^{l}_{i},y_{i})=(\bar{z}_{k}^{l},\bar{y}_{k})$ for some $k$. Then, by Proposition˜3.1 and the subsequent remark (5), any subsequent layer, $W_{j}^{*}$ with $j>l$, has rank at most $K$.
@@ 3.1 Clustered Data and Total Cluster Variation
Assume $\mathcal{C}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{K}\}$ to be a $K$-partition of the input data $\mathcal{X}$ and corresponding labels $\mathcal{Y}$, i.e., for any $(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$ there exists a unique $\mathcal{C}_{k}\in\mathcal{C}$ such that $(x_{i},y_{i})\in\mathcal{C}_{k}$. We denote by $N_{k}=|\mathcal{C}_{k}|$ the cardinality of the $k$-th family and adopt the notation
$\mathcal{X}\times\mathcal{Y}=\bigcup_{k=1}^{K}\bigcup_{i\in\mathcal{C}_{k}}\{(x_{i},y_{i})\}=\bigcup_{k=1}^{K}\bigcup_{i=1}^{N_{k}}\{(x_{i}^{k},y_{i}^{k})\},$
where with a small abuse of notation we use $\mathcal{C}_{k}$ to denote also the set of indices whose corresponding data points are contained in $\mathcal{C}_{k}$. Now let $\bar{x}_{k}=\bar{z}^{0}_{k}$, $\bar{y}_{k}$ and $\bar{z}^{l}_{k}$ be the centroids of the sets $\{x_{i}\}_{i\in\mathcal{C}_{k}}$, $\{y_{i}\}_{i\in\mathcal{C}_{k}}$ and $\{z^{l}_{i}\}_{i\in\mathcal{C}_{k}}$, respectively, i.e.,
@@
This quantity measures the within-cluster variance (see, e.g., hastie2009elements ) of the data at layer $l$ with respect to a prescribed partition and the corresponding centroids.
@@
Note that, for $K$ sufficiently smaller than $N$, a small value of $\operatorname{\mathrm{TCV}}_{K,l}$ indicates an intermediate collapse of the network’s features at layer $l$. Moreover, being $\operatorname{\mathrm{TCV}}_{K,l}$ defined through a variational problem to get the tightest bound possible, it is trivially upper bounded by the total cluster variation on the classes. Therefore, neural collapse in the classical sense implies $\operatorname{\mathrm{TCV}}_{K,l}=0$.

### v3

@@ Abstract
We present a unified theoretical framework connecting the first property of Deep Neural Collapse (DNC1) to the emergence of implicit low-rank bias in nonlinear networks trained with $L^{2}$ weight decay regularization. Our main contributions are threefold. First, we derive a quantitative relation between the Total Cluster Variation (TCV) of intermediate embeddings and the numerical rank of stationary weight matrices. In particular, we establish that, at any critical point, the distance from a weight matrix to the set of rank-$K$ matrices is bounded by a constant times the TCV of earlier-layer features, scaled inversely with the weight-decay parameter. Second, we prove global optimality of DNC1 in a constrained representation-cost setting for both feedforward and residual architectures, showing that zero TCV across intermediate layers minimizes the representation cost under natural architectural constraints. Third, we establish a benign landscape property: for almost every interpolating initialization there exists a continuous, loss-decreasing path from the initialization to a globally optimal, DNC1-satisfying configuration. Our theoretical claims are validated empirically; numerical experiments confirm the predicted relations among TCV, singular-value structure, and weight decay. These results indicate that neural collapse and low-rank bias are intimately linked phenomena arising from the optimization geometry induced by weight decay.
1 Introduction
@@
Recent years have seen growing theoretical and experimental evidence of low-rank properties of deep neural networks. For example, it has been noted that a large portion of singular values from linear layers in different models can be removed almost without affecting performance, both for feedforward and residual networks (Schotthöfer et al., 2022; Idelbayev and Carreira-Perpiñán, 2020). Analogous results have been shown on feedforward and convolutional networks for image classification (Li et al., 2019; Wang et al., 2021; Khodak et al., 2021; Schotthöfer et al., 2022) as well as on Recurrent Neural Networks (RNNs) trained to learn neural dynamics in neuroscience tasks (Schuessler et al., 2020; Pellegrino et al., 2023).
@@
Alongside the wealth of experimental evidence of low-rank properties of neural networks, several recent works have investigated the low-rank bias phenomenon from the theoretical point of view, providing evidence of decaying singular values in deep linear architectures (Arora et al., 2019; Feng et al., 2022; Huh et al., 2022) and of low-rank properties of networks’ Hessian maps (Singh et al., 2021), analyzing low-rank geometric properties of training gradient flow dynamics for deep linear networks (Bah et al., 2022), showing that low-rank bias is related to weight decay (Galanti et al., 2023; Súkeník et al., 2024a; Kuzborskij and Yadkori, 2025), and to training invariances (Le and Jegelka, 2022; Ziyin et al., 2025). In particular, the authors relate the rank of the weight matrices during SGD to the batch size adopted during the training of the network. However, based only on the batch size, the results from (Galanti et al., 2023) become uninformative when the batch size is comparable to the number of parameters of the neural network.
@@
A related phenomenon commonly observed during the training of classifiers is “neural collapse” (Papyan et al., 2020; Zhu et al., 2021) and more generally "deep (or intermediate) neural collapse" Parker et al. (2023); Rangamani et al. (2023b); Súkeník et al. (2024b; a). This phenomenon is rigorously described in the literature using four properties named NC1–4, the first of which (NC1) essentially states that the within-class variability of intermediate representations goes to zero even for layers in the middle of a deep neural network (see Definition˜3.3). The optimality of neural collapse at the last layer has been extensively investigated and proved in different settings, otherwise only few extensions have been proposed for its deep version.
In particular, the emergence of the neural collapse phenomenon at the last layer was first empirically observed in (Papyan et al., 2020), and proven to be optimal in different settings, such as Unconstrained Feature Model (UFM) (Fang et al., 2021; Mixon et al., 2020; Zhu et al., 2021), with constrained UFM (Dang et al., 2024; Yaras et al., 2023), with cross-entropy loss (E and Wojtowytsch, 2021; Kunin et al., 2023; Hong and Ling, 2024), Neural Tangent Kernel (NTK) block structure aligned with class labels (Seleznova et al., 2023), linear tail and wide architecture (Jacot et al., 2025), effects of the loss function (Zhou et al., 2022b), and in the context of Graph Neural Networks (GNNs) (Kothapalli et al., 2023). Moreover, some works analyzed the dynamics leading to neural collapse, from the observation of a central path dynamics (Han et al., 2022), to the use of large network dynamics (Jacot et al., 2025). Finally, some authors characterize stability of the stationary points of UFM in the regularized setting (Ji et al., 2022; Zhu et al., 2021; Zhou et al., 2022a), and in the Riemannian setting, with feature and weights constrained to oblique manifolds (Yaras et al., 2023).
The analysis of the neural collapse phenomenon beyond the last layer was first empirically studied by (Rangamani et al., 2023b), and addressed under the name of “deep” (or “intermediate") neural collapse phenomenon (DNC), where notably the NC1 property is observed to hold even at intermediate layers. Proving $\operatorname{\mathrm{DNC1}}$ optimality has proven more challenging than NC because it requires switching from the linear setting of the last layer to the nonlinear one of the intermediate layers. A theoretical analysis of this phenomenon has been conducted in (Súkeník et al., 2024b; a) that establish $\operatorname{\mathrm{DNC1}}$ optimality in the Deep-UFM model but restrict to binary classification or require specific architectural assumptions (L-DUFM, pyramidal widths). In particular, the key limitation of prior work about $\operatorname{\mathrm{DNC1}}$ is that it studies only simplified linear/UFM models, not able to take into account the full extent of architectural constraints.
1.1 Main Contributions
@@
In this work we show, for the first time, a rigorous connection between the phenomenon of deep neural collapse and low-rank implicit bias for general nonlinear networks. The main contribution is theoretical and, in particular, our main results, which hold in the presence of weight decay, can be summarized as follows:
@@
  • in Section˜3 we prove that within-class variability of hidden layers during training (including the initial data set) has a quantifiable effect on the numerical rank of the network parameters as the latter decays with the total cluster variation of the hidden representations (Theorem˜3.5);
@@
  • in Section˜4 we show that zero total cluster variation for all intermediate layers is optimal in a fairly general setting, both for feedforward and residual networks (Theorem˜4.2);
  • Theorem˜4.4 establishes that for (almost) every interpolating initial condition, there exists a loss-decreasing trajectory connecting the initial condition to a collapsed configuration, thereby offering an explanation to the consistent empirical observation of deep neural collapse (Rangamani et al., 2023b).
@@
Overall, the combination of the results above proves that in the regime of small weight decay regularization, zero within-class variability for all intermediate representations is globally optimal under some natural constraints, therefore showing that global minima in parameter space are low-rank matrices. Moreover, convergence to these highly structured solutions may be regarded as universal in the sense that, there is essentially no loss barrier between interpolating features to deep-neural collapsed ones, which constitutes a step forward towards explaining the universality in the emergence of (deep) neural collapse.
Our paper improves upon the results presented in the related literature in different aspects. First, we show that the rank of stationary points in parameter space is controlled by $\operatorname{\mathrm{DNC1}}$, a metric defining the first property of neural collapse. In particular, our result also holds for stationary points that do not satisfy DNC2/DNC3. Notably, the results in (Súkeník et al., 2024a) suggest that there exist settings where our results become crucial, indeed the authors show that, in such settings, while $\operatorname{\mathrm{DNC1}}$ keeps being optimal, the other properties may fail.
Second, the setting in which we prove our results is not a simplified model, but an exact representation of a feedforward/ResNet architecture with two additional constraints. The analysis is based on the extrinsic view of a neural network architecture as a constraint in the space of representable intermediate embeddings (Jacot et al., 2022; Han et al., 2022). Finally, in this generalized setting, we prove a benign property of the loss landscape, which could explain the frequent emergence of neural collapse in intermediate representations as observed in (Rangamani et al., 2023b).
The remainder of the paper is organized as follows. In Section˜2 we introduce notation and set up the optimization framework. In Section˜3 we prove the rank–TCV bound and discusses its implications. Section˜4 contains the optimality and landscape results, and Section˜5 validates the theory empirically.
@@ 2 Setting and Preliminaries
to learn from the given data and labels. In equation equation 1, $\sigma_{l}$ are scalar nonlinear activation functions acting entrywise and $\mathit{\Theta}$ denotes all trainable parameters $\mathit{\Theta}=(W_{1},\dots,W_{L},b_{1},\dots,b_{L})$. We write
@@
to denote the training parameters and concatenation of layers from $l_{1}+1$ to $l_{2}$, respectively. We adopt the convention that, when we do not specify the lower index then it starts from zero, and when we omit the top one it ends at the overall depth $L$, namely, $\mathit{\Theta}^{l_{2}}:=\mathit{\Theta}_{0}^{l_{2}}$, $f^{l_{2}}:=f^{l_{2}}_{0}$ and $\mathit{\Theta}_{l_{1}}:=\mathit{\Theta}_{l_{1}}^{L}$, $f_{l_{1}}:=f^{L}_{l_{1}}$. Notice that, with this notation, $\mathit{\Theta}=\mathit{\Theta}_{0}=\mathit{\Theta}^{L}=\mathit{\Theta}_{0}^{L}$, $f=f_{0}=f^{L}=f_{0}^{L}$ and $\mathit{\Theta}^{0}$ is empty. Finally, for the sake of brevity, we write $\mathcal{Z}^{l}:=\{z^{l}_{i}\}_{x_{i}\in\mathcal{X}}$ to denote the set of outputs of the $l$-th layer, where $z^{l}_{i}=f^{l}(\mathit{\Theta}^{l},x_{i})$ is defined as in equation 1 with input data $x_{i}$, and $\mathcal{Z}^{0}=\mathcal{X}$.
@@
3 A Link Between Low-Rank and Deep-Neural Collapse
@@
The intuition is that Equation equation 4 resembles an “eigenvalue equation” for the nonlinear operator $\nabla\operatorname{\mathcal{L}}_{0}$. In other words, for a fixed weight decay parameter $\lambda$, training the neural network corresponds to finding a nonlinear eigenvector $\mathit{\Theta}^{*}$ having $\lambda$ as eigenvalue. In addition, equation equation 4 immediately implies that whenever the operators $\nabla_{W_{l}}\operatorname{\mathcal{L}}_{0}$ have low rank (i.e., they take values in some low-rank space), then also any weight matrix $W^{*}$ composing the eigenvector $\mathit{\Theta}^{*}$ has necessarily low rank. In the following, we investigate this relationship more in depth, providing sufficient conditions to have approximately low-rank gradients.
@@
Let $f(\mathit{\Theta},x)$ be a feedforward neural network as in equation 1 and $\operatorname{\ell}:\mathbb{R}^{c}\times\mathbb{R}^{c}\to\mathbb{R}$ be a differentiable loss function as in equation 3. Then, for any $\{(x_{i},y_{i})\}_{i=1}^{K}\subseteq\mathbb{R}^{d}\times\mathbb{R}^{c}$, we have $\operatorname{rank}\Big(\operatorname{\nabla}_{W_{l}}\sum_{i=1}^{K}\operatorname{\ell}\big(f(\mathit{\Theta},x_{i}),y_{i}\big)\Big)\leq K$.
@@
In addition, Proposition˜3.1 immediately yields an example where gradients have necessarily low rank. Consider the case of degenerate data, i.e., assume that there exist $K$ pairs $\{(\bar{x}_{k},\bar{y}_{k})\}_{k=1}^{K}$ such that any data point $(x_{i},y_{i})$ is equal to some $(\bar{x}_{k},\bar{y}_{k})$. Then, by Proposition˜3.1, $\operatorname{\nabla}_{W_{l}}\sum_{i}\operatorname{\ell}(\mathit{\Theta},x_{i},y_{i})$ has rank at most $K$. But the right-hand side of equation 4 has rank at most $K$, hence any critical point of the loss function $\operatorname{\mathcal{L}}_{\lambda}$ necessarily corresponds to a neural network having all layers of rank at most $K$. The same argument can be extended to the situation of total intermediate neural collapse (Rangamani et al., 2023a), where the intermediate hidden set $\mathcal{Z}^{l}$ consists of at most $K$ distinct points. Indeed, assume that for a stationary point $\mathit{\Theta}^{*}$ there exists a layer $l$ and points $\{(\bar{z}_{k}^{l},\bar{y}_{k})\}_{k=1}^{K}$ such that any pair $(z^{l}_{i},y_{i})=(\bar{z}_{k}^{l},\bar{y}_{k})$ for some $k$. Then, by Proposition˜3.1 and the subsequent remark equation 5, any subsequent layer, $W_{j}^{*}$ with $j>l$, has rank at most $K$.
@@ 3.1 Clustered Data and Total Cluster Variation
Assume $\mathcal{C}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{K}\}$ to be a $K$-partition of the input data $\mathcal{X}$ and corresponding labels $\mathcal{Y}$, i.e., for any $(x_{i},y_{i})\in(\mathcal{X},\mathcal{Y})$ there exists a unique $\mathcal{C}_{k}\in\mathcal{C}$ such that $(x_{i},y_{i})\in\mathcal{C}_{k}$. We denote by $N_{k}=|\mathcal{C}_{k}|$ the cardinality of the $k$-th family and adopt the notation $(\mathcal{X},\mathcal{Y})=\bigcup_{k=1}^{K}\bigcup_{i\in\mathcal{C}_{k}}\{(x_{i},y_{i})\}=\bigcup_{k=1}^{K}\bigcup_{i=1}^{N_{k}}\{(x_{i}^{k},y_{i}^{k})\}$, where with a small abuse of notation we use $\mathcal{C}_{k}$ to denote also the set of indices whose corresponding data points are contained in $\mathcal{C}_{k}$. Now let $\bar{x}_{k}=\bar{z}^{0}_{k}$, $\bar{y}_{k}$ and $\bar{z}^{l}_{k}$ be the centroids of the sets $\{x_{i}\}_{i\in\mathcal{C}_{k}}$, $\{y_{i}\}_{i\in\mathcal{C}_{k}}$ and $\{z^{l}_{i}\}_{i\in\mathcal{C}_{k}}$, respectively, i.e.,
@@
This quantity measures the within-cluster variance (see, e.g., (Hastie et al., 2009)) of the data at layer $l$ with respect to a prescribed partition and the corresponding centroids.
@@
Note that, for $K$ sufficiently smaller than $N$, a small value of $\operatorname{\mathrm{TCV}}_{K,l}$ indicates an intermediate collapse of the network’s features at layer $l$. Moreover, being $\operatorname{\mathrm{TCV}}_{K,l}$ defined through a variational problem to get the tightest bound possible, it is trivially upper bounded by the total cluster variation on the classes. These considerations are directly connected with the notion of deep (or intermediate) neural collapse (Súkeník et al., 2024b; Papyan, 2020; Rangamani et al., 2023b), whose first characterizing property is recalled below.
Definition 3.3 (Deep Neural Collapse 1).
We say that a neural network $f$ trained on a dataset $(\mathcal{X},\mathcal{Y})$ satisfies $\operatorname{\mathrm{DNC1}}$ at layer $l$ if $\operatorname{\mathrm{TCV}}_{K,l}=0$ for $K=c$ and the minimizer in equation 7 is given by the partition corresponding to the classes.